In 1912 the U.S. government began issuing licenses to radio stations. Each station was given a unique three-letter "call sign." By
international agreement, the United States received rights to all call signs beginning with the letters W, N, and K. Radio stations
in the western United States were given call signs starting with K. Stations in the East were given call signs starting with W. (N
was reserved for use by the U.S. Navy.)
(a) How many three-letter call signs (like WGO) start with the letter W?
(b) How many three-letter call signs starting with W or K were available for U.S radio stations?

Answer by ikleyn(52879)

.

I assume that 26 capital letters of English alphabet are used/allowed for each of the three positions.


(a)  First of the three positions is letter W.

     Two other positions bring 26 options each, independently.

     So, 26*26 = 676 is the ANSWER for (a).



(b)  Use the same logic.

     26*26 = 676 call signs start with W and 26*26 = 676 other call signs start with K.

     In all, 676 + 676 = 1352 is the ANSWER for (b).

Solved.